OpenStudy (anonymous):

What is the oxidation number of each phosphorus atom (P) in the compound calcium phosphite, Ca3(PO3)2? -2 -3 +3 +6

OpenStudy (chmvijay):

(PO3)2 ^-6 is having -6 charge single (PO3)-3 will have -3 charge P+3(-2) = -3 P-6=-3 P=-3+6=+3

OpenStudy (anonymous):

Tyyy smmm!